a gambler finishes a session of black jack with $5 chips and $25 chips. If he has 45 chips in all, with a total valie of $465, how many of each kind of chip does the gambler have?
Answer by jorel1380(3719)

x+y=45
5x+25y=465
5x+5y=225
20y=240
y=12
x=33
He has 33 $5 chips and 12 $25 chips..